OpenStudy (anonymous):

Glenn rolls a 6 sided die. what is the probability that he gets a number greater than 2 or an odd number?

OpenStudy (anonymous):

Number greater than 2= 4/6 Odd number= 3/6

OpenStudy (anonymous):

so 12/36 ?

OpenStudy (colouringme):

when a question using the word "..and..." multipy every probabilities that are required. when it used the word "...or..." add them. In this case, he has to get greater than two OR odd . which means, you have to multiply the probabilities ; \[(\frac{ 4 }{ 6 } + \frac{ 3 }{ 6 })\] which I think the answer is 1, because you will always get odd OR more than two. And you can only use those two rules above if the dice is rolled twice, or there's two dice rolled together

OpenStudy (anonymous):

im confused lol

OpenStudy (colouringme):

you must get the result odd : 1, 3,5 or greater than 2 : 3,4,5, 6 so, you have ... options out of 6 probability of the number that will show up...

OpenStudy (colouringme):

does it help? Ignore my first reply :x

OpenStudy (anonymous):

Yes haha thank you!

OpenStudy (colouringme):

so the answer is.... :)

OpenStudy (anonymous):

5/6 ?

OpenStudy (colouringme):

yeah! :D

OpenStudy (anonymous):

thank you so much love (:

OpenStudy (colouringme):

No problem :)